# Calibrated and uncertain? Evaluating uncertainty estimates in binary classification models

### v1

@@ 5.1 Calibration results
First we attempt to answer Q1 by evaluating if the estimated probabilities fluctuate around the LRFD and approach it as we increase the amount of training data. Figure 3 and Figure 4 shows the estimated probabilities $\bar{Q}(c^{2}=1|\mathbf{x})$ and uncertainties $u_{Q}(c^{2}=1|\mathbf{x})$ as a function of polar coordinate radius $r=|\mathbf{x}|$ with error bars indicating the entire spread of estimates over polar angle $\phi$, calculated on the test set of each of the six algorithms from training on 500 and 5000 training data points from dataset A and B respectively. The figures show that probability estimates are well calibrated for all algorithms from around $0\leq|\mathbf{x}|\leq 30$ for dataset A and $0\leq|\mathbf{x}|\leq 20$ for dataset B. As we add more training data, the estimates concentrate around the LRFD and the error bars shrink. From around $|\mathbf{x}|\geq 30$ and $|\mathbf{x}|\geq 20$ for dataset A and B respectively, some of the estimates do not fluctuate around the LRFD and are no longer calibrated, which becomes more apparent as we move further away from the bulk of the training data. The behavior in the tails is further covered in the results of Q3.
@@
Figure 5 shows the five different calibration-related summary statistics as well as the global accuracy calculated on the test set as a function of increasing number of training data points. The plots show that all models converge towards the LRFD as the data size increases, but the metrics are quite noisy and do not always give the same relative ranking to each model. The neural networks are trained specifically to minimize the cross entropy loss and a decrease in this metric does not always correspond to a decrease in for example estimated calibration error or even KL divergence. For dataset A, the GP model scores are better relative to many of the other models, as the LRFD of that dataset aligns better with the prior (bias) of the GP. For dataset B the opposite is true. These plots also illustrate that estimated calibration error and cross entropy loss, which are evaluated using only the labels of the test set and not the LRFD can be fairly good estimates of the usually intractable Wasserstein-1 distance and mean KL-divergence metrics (if the test set is large enough).
@@
Figure 6: Top: Average estimated out-of-distribution probabilities (left), out-of-distribution uncertainty (middle) and test set uncertainty (right) as a function of $N_{\text{train}}$ for dataset A (top row) and B (bottom row). The error bars indicate the entire spread of the data over polar angle $\phi$. The colors are similar to Figure 5. The dashed black line in the middle plot indicates the standard deviation of the posterior of a Bernoulli distribution with one data point of class 2 with prior $\text{Beta}(1,1)$.
5.2 Data-size dependency for uncertainty results
@@
Next we investigate Q2 by checking if the uncertainties for both in-distribution and OOD data goes down as we increase the training data size. Figure 6 shows the average estimated probabilities and uncertainties for the OOD test points with $700<|\mathbf{x}|<1000$ as well as average uncertainties for the in-distribution test set as a function of increasing number of training data points. For the OOD test points, all algorithms except for GP and DPMM produce approximately constant uncertainties. The standard deviation of the posterior of a Bernoulli distribution with one data point of class 2 with prior $\text{Beta}(1,1)$, as described by Equation 3 is also plotted for comparison. GP uncertainties decrease, which is due to the calculated approximate optimal value for the output variance of the Radial Basis Kernel decreasing with increasing nr of data points (see Appendix A). The DPMM uncertainties may decrease due to numerical errors because of increasingly small probability density estimates in the tail, as expanded on in the next results section. For all algorithms except EDL, the uncertainties on the test set tend to go down as we increase the size of the training set. For EDL however, the uncertainties increase and for dataset B, CL average uncertainties stay almost constant as we increase the data size.
5.3 Tail estimates and out-of-distribution probability and uncertainty results
@@
Finally we answer Q3 by checking if uncertainty estimates increase as we move away from the bulk of the training data. We go back to examining Figure 3 and Figure 4, with a focus on the tails, defined loosely to be $|\mathbf{x}|\geq 30$ and $|\mathbf{x}|\geq 20$ for dataset A and B respectively. For dataset A, the neural network-based algorithms tend to first underestimate the class 2 probabilities, then overestimate them. For dataset B, they almost perfectly align with the LRFD. The nonparametric Bayesian inference algorithms on the other hand underestimate the probabilities in the tails relative to the LRFD for both datasets, giving estimates closer to the noninformative prior probability estimate of 0.5. The spread over polar angles of the estimates (the height of the error bars in the plot) as a function of increasing amounts of data in this data range also differ between the neural network and the nonparametric Bayesian models. The neural network models have error bars that shrink with increasing amounts of data regardless of whether the estimates are close or not to the LRFD, while the nonparametric Bayesian models have error bars that do not shrink (the DPMM error bars even grow) for test points in the tails. The uncertainty estimates of all algorithms approximately follow the estimated probabilities in the calibrated areas as the standard deviation of the estimated frequency of a Bernoulli distribution is proportional to $\sqrt{\nu(1-\nu)}$, which results in a local maxima around $\nu=0.5$. The height of this maxima is different for all the models, indicating that the uncertainty values are highly model dependent. In the tails of the test set distribution, uncertainties increase for the nonparametric Bayesian models. The neural network models have flat or decreasing uncertainties in the tail.
@@
The first two panels of Figure 6 show the average estimated probabilities and uncertainties on the OOD datasets. For most deep learning models, the probability estimates are approximately 1. The only exception is EDL which for dataset A has very noisy estimates centered around 0.9. The NN-algorithms all also have estimated uncertainties of approximately 0. This is the correct value for a variance proportional to $\sqrt{\nu(1-\nu)}$ when $\nu$ is 0 or 1, and so does not indicate that there is some fundamental error in the approximation of the variance of the distribution. The exception is the NNE uncertainty estimates for dataset A $N_{\text{train}}=5000$ which are non zero and extremely noisy. This anomaly corresponds to diverging estimates in the ensemble, which seems to be a rather unusual result when the ensemble is trained well. The NN models therefore do not show a strong, consistent increase in uncertainty with decreasing density $p(\mathbf{x})$. The GP algorithm produces probability estimates of approximately 0.5 and high uncertainty for OOD data points for both datasets independent of data size. The DPMM algorithm also produces high uncertainty for OOD data and OOD probability estimates close to 0.5 for the smallest amount of training data, but the probability estimates diverge from 0.5 as we increase the number of data points, pushing the estimated distribution towards 1. This is unexpected as the expected out-of-distribution frequency of a Bernoulli distribution using the Bayes-Laplace (uniform) prior as our DPMM algorithm does, should give an estimate of 0.5. To calculate the conditional distribution however, the algorithm actually approximates the joint distribution, which gives very small probabilities for the data in low density areas. The results may therefore be influenced by biased numerical errors or possibly pathologies of the underlying mathematical framework (Dunson and Bhattacharya, 2011).
@@
To study the effect of hyperparameter choice on our deep learning results we have performed a grid search over relevant hyperparameters such as the size of the neural networks, the learning rate, the weight decay and some algorithm specific hyperparameters. Detailed results are presented in Appendix B. We find that some settings of hyperparameters give rise to non-extreme OOD probability estimates, but these correspond to networks that have not learned well and perform badly on the in-distribution test set. The majority of the trained networks result in extreme probability estimates similar to the ones presented in the main part of the study. Some of the hyperparameter choices also give rise to higher OOD uncertainty than in the main findings. These are also connected to networks that perform worse and often significantly so. All in all, our grid search corroborates our main findings and provide some interesting additional information on the behavior of the algorithms.
As an additional test we performed the same experiments with a dataset with two identically distributed classes. Predictions on OOD data from NNE, CL and MCD varied greatly for different polar angles. For small learning rates, EDL estimated probabilities close to 0.5 and uncertainty close to 0. GPs also estimated probabilities close to 0.5 and uncertainty close to 0, while DPMMs estimated probabilities close to 0.5 and uncertainties close to 0.3. For more information see Appendix C.
@@ 6 Discussion and conclusion
  • the average uncertainty on the test set go down as we increase the number of training data points for all algorithms except evidential deep learning which go up, indicating that something might be wrong in assumptions or implementation;
@@
  • the out-of-distribution average uncertainty as a function of number of training data points is high and goes down for the Gaussian process classifier and Dirichlet process mixture model, and is low or zero and constant for the deep learning models;
@@
  • for out-of-distribution data all deep learning algorithms estimated extreme probabilities, while the nonparametric Bayesian models estimate probabilities around 0.5, with the Dirichlet process mixture model estimating slightly higher probabilities as the number of training data points increase, indicating some numerical instability or other error in implementation.
@@
The DPMM is explicitly defined as having a $\text{Beta}(1,1)$ prior for class probability, and results support the qualitative correctness of the approximate posterior. For the neural networks and GP, the prior is implicitly defined, and so it is harder to evaluate the cohesiveness of the results. It is possible to interpret the deep learning results as representative of approximate posterior distributions with $\text{Beta}(0,0)$ priors for class probability. This suggests that changing the implicit prior for the deep learning models through modifying the loss function could change the observed behavior. However, it might not be a good idea to assign those kind of interpretations post-hoc, as we might put too much trust in a method without a principled theoretical foundation. Instead, we might consider if the approximations introduced by variational inference are the cause of the deviation from desired results.
@@
Although neural networks with rectified linear (ReLU) activation functions can in theory approximate any nonlinear function of the data locally, they eventually become simple linear functions of the input features if one moves in any direction far enough (Hein et al., 2019). This property of piecewise linear neural networks, combined with using the softmax function for normalization of the last layer of the network, gives rise to class probabilities that asymptotically approach extreme values of zero or one. Therefore the estimated extreme probabilities for out-of-distribution data may simply be a feature of using ReLU neural networks as the family of parameterized test distributions in variational inference. Similar to more well known parametric methods such as linear regression we should not expect our model to be informative or correct if we extrapolate too far. Changing the architecture and activation functions to better reflect the kind of statistical model we seek is an obvious step to mitigate the problem of model misspecification.
@@
This study only investigates data which is sampled from a relatively simple two dimensional distribution and therefore does not represent the complexities present in real-world data. It is noteworthy that even for this relatively simple problem, many of the learning algorithms do not produce uncertainties which can be used by scientists to decide if they should trust the estimates. Although the uncertainty estimates produced by the deep learning algorithms presented in this study were not especially useful for this type of data, they might work better for data with for example discrete or bounded distributions, or for data with statistically independent input features. We strongly recommended that scientists create their own toy data that more closely matches the mathematical properties of their own data of interest to see how the different machine learning algorithms can be expected to behave in their case.
Regardless of the mathematical reason for the observed behavior and the inherent limitation of empirical studies, we can take away some key lessons from the experiment. Deep learning-based class probability estimates and uncertainties do not necessarily exhibit the behavior we expect or desire for out-of-distribution data. This may be partly due to a lack of understanding of what can reasonably be expected from a statistical method, like the fact that it is not possible to have unbiased frequency estimates which also extrapolate to a specific value in the $N\rightarrow 0$ limit. It can also be because we do not yet have a clear understanding of the implicit statistical models we create or the inherent limitations of the approximations we make in our inference algorithms to facilitate faster learning and compute intractable integrals. In any case, this study should serve as a reminder that purely data-driven complex statistical models and learning algorithms should not be trusted out of the box. Sanity checks with intuitive toy-problems such as this one are a useful tool to explore the way these algorithms behave, and we hope this study can serve as an inspiration and a guide for future machine learning development work in the natural sciences.

### v2

@@ 5.1 Q1: Calibration
Figure 3 and Figure 4 shows the estimated probabilities $\bar{Q}(c^{1}=1|\mathbf{x})$ and uncertainties $u_{Q}(c^{1}=1|\mathbf{x})$ as a function of polar coordinate radius $r=|\mathbf{x}|$ with error bars indicating the entire spread of estimates over polar angle $\phi$, calculated on the test set of each of the six algorithms from training on 500 and 5000 training data points from dataset A and B respectively. The figures show that probability estimates are reasonably well calibrated for all algorithms from around $0\leq|\mathbf{x}|\leq 30$ for dataset A and $0\leq|\mathbf{x}|\leq 20$ for dataset B. As we add more training data, the estimates concentrate around the LRFD and the error bars shrink. From around $|\mathbf{x}|\geq 30$ and $|\mathbf{x}|\geq 20$ for dataset A and B respectively, some of the estimates do not fluctuate around the LRFD and are no longer calibrated, which becomes more apparent as we move further away from the bulk of the training data. The behavior in the tails is further covered in the results of Q3.
@@
Figure 5 and 6 shows the five different calibration-related summary statistics as well as the global accuracy calculated on the test set and the data points in the tails of the test set as a function of increasing number of training data points. The plots show that all models converge towards the LRFD and the empirical estimates of the LRFD as the data size increases, but the metrics are quite noisy and do not always give the same relative ranking to each model. The neural networks are trained specifically to minimize the cross entropy loss and a decrease in this metric does not always correspond to a decrease in the other metrics. For dataset A, the GP model scores for the entire test set are better relative to many of the other models. For dataset B the opposite is true. Panel b) in Figure 6 shows how calibration scores get worse in the tails for the conservative estimates of the GP and DPMM models. However, this is not necessarily reflected in the scores averaged over the entire dataset as there are naturally much fewer data points in the tails. Reliability diagrams are presented in Appendix D and show that the algorithms, when given enough data, are all well calibrated for bins over $\bar{Q}(c^{1}=1|\mathbf{x})\in[0,1]$ which do not just include data points in the tails.
(a) Dataset A
(b) Dataset B
@@
Figure 7: Top: Average estimated out-of-distribution probabilities (top rows) and uncertainties (bottom rows) as a function of $N_{\text{train}}$ for the test set (left), the OOD-near test set (middle) and the OOD-far test set (right). The error bars indicate the entire spread of the data over polar angle $\phi$. The colors are similar to Figure 5. The dashed black line in the plots in the bottom rows indicates the standard deviation of the predictive posterior of a binomial model with $N=0$ and prior $\text{Beta}(1,1)$.
5.2 Q2: Uncertainty vs number of training points
Next we investigate Q2 by checking if the class probability uncertainty estimates go down as we increase the training data size. The results indicate that the average uncertainty evaluated over the test set is either constant or goes down for all models except EDL which exhibits an increase in uncertainty. For the OOD data we see the same trends, but on the OOD-near test set for dataset A we also see an increase in uncertainty for DPMM.
@@ 5.2 Q2: Uncertainty vs number of training points
Figure 7 shows the average estimated probabilities and uncertainties for the test set, the OOD-near test set ($80\leq r\leq 120$) and the OOD-far test set ($700\leq r\leq 1000$). For all algorithms except EDL, the uncertainties on the test set tend to go down or remain constant as we increase the size of the training set. For EDL however, the uncertainties increase. For the OOD test points, all deep learning algorithms except EDL produce approximately constant uncertainties with value 0. The EDL uncertainties increase as a function of $N_{\text{train}}$ for the OOD-near test set for dataset A, and are approximately zero and constant for the rest. GP uncertainties decrease, which is due to the calculated approximate optimal value for the output variance of the Radial Basis Kernel decreasing with increasing nr of data points (see Appendix A). DPMM uncertainties decrease as a function of $N_{\text{train}}$ for the test sets and for the OOD test sets for dataset B, but increase for the OOD test sets for dataset A. The DPMM uncertainties may exhibit this increase due to numerical errors because of increasingly small probability density estimates in the tail, as expanded on in the next results section.
5.3 Q3: OOD uncertainty
Finally we answer Q3 by checking if uncertainty estimates increase as we move away from the bulk of the training data. Our main findings are that for the deep learning models using the ReLU activation function, uncertainties are not higher in the tails of the test set or in the OOD-near and OOD-far datasets. For the nonparametric Bayesian models, uncertainties were higher for OOD data, but the DPMM shows some non-intuitive behavior. In an additional experiment using completely identical data distributions, only the DPMM produced probability and uncertainty estimates that had high uncertainty for OOD data.
@@ 5.3 Q3: OOD uncertainty
We go back to examining Figure 3 and Figure 4, with a focus on the tails, defined loosely to be $|\mathbf{x}|\geq 30$ and $|\mathbf{x}|\geq 20$ for dataset A and B respectively. For dataset A, the neural network-based algorithms tend to first underestimate the class 1 probabilities, then overestimate them. For dataset B, they almost perfectly align with the LRFD. The nonparametric Bayesian inference algorithms on the other hand underestimate the probabilities in the tails relative to the LRFD for both datasets, giving class probability estimates closer to the noninformative prior probability estimate of 0.5. In the tails of the test set distribution, class probability uncertainties increase for the nonparametric Bayesian models. The neural network models have flat or decreasing uncertainties in the tail. The uncertainty estimates of all algorithms approximately follow the relation $\sigma_{\nu}\propto\sqrt{\bar{\nu}(1-\bar{\nu})}$, which results in a local maxima around $\nu=0.5$. The height of this maxima is different for all the models, indicating that the uncertainty values are highly model dependent.
@@
The first two panels of Figure 7 show the average estimated probabilities and uncertainties evaluated over the OOD datasets. For most deep learning models, the probability estimates are approximately 1. The only exception is EDL which for dataset A has very noisy estimates centered around 0.9. The NN-algorithms all also have estimated uncertainties of approximately 0. This is the correct value for a standard deviation proportional to $\sqrt{\nu(1-\nu)}$ when $\nu$ is 0 or 1, and so does not indicate that there is some fundamental error in the approximation of the variance of the distribution. The exception is the NNE uncertainty estimates for dataset A $N_{\text{train}}=5000$ which are non zero and extremely noisy. This anomaly corresponds to diverging estimates in the ensemble, which seems to be a rather unusual result when the ensemble is trained well. The NN models therefore do not show a strong, consistent increase in uncertainty with decreasing density $p(\mathbf{x})$. The GP algorithm produces probability estimates of approximately 0.5 and high uncertainty for OOD data points for both datasets independent of data size. The DPMM algorithm also produces high uncertainty for OOD data and OOD probability estimates close to 0.5 for the smallest amount of training data, but the probability estimates diverge from 0.5 as we increase the number of data points, pushing the estimated distribution towards 1. This is not what we would expect to see as the DPMM uses a Bernoulli distribution with a Bayes-Laplace (uniform) prior, which should give posterior class probabilities of 0.5 where there is no data. To calculate the conditional distribution however, the algorithm actually approximates the joint distribution, which gives very small probabilities for the data in low density areas. The results may therefore be influenced by biased numerical errors or possibly pathologies of the underlying mathematical framework (Dunson and Bhattacharya, 2011). The inspected MCMC diagnostics did not indicate any problems, and the behavior persisted if we changed the hyperparameter of the Beta-prior, indicating that this is a stable numerical issue of the implementation.
@@
To study the effect of hyperparameter choice on our deep learning results we have performed a grid search over relevant hyperparameters such as the size of the neural networks, the learning rate, the weight decay and some algorithm specific hyperparameters. Detailed results are presented in Appendix B. We find that in ReLU neural networks some settings of hyperparameters give rise to non-extreme OOD probability estimates, but these correspond to networks that have not learned well and perform badly on the in-distribution test set. The majority of the trained networks result in extreme probability estimates similar to the ones presented in the main part of the study. Some of the hyperparameter choices also give rise to higher OOD uncertainty than in the main findings. These are also connected to networks that perform worse and often significantly so. We also tested changing the activation function in the neural networks to tanh, with the results presented in Appendix C. In this case, the NNE, CL and EDL algorithms sometimes produced uncertainties that increased for OOD data without sacrificing accuracy. However, the behavior was not consistent across the datasets or number of training data points. All in all, our grid search corroborates our main findings and provide some interesting additional information on the behavior of the algorithms.
(a) Neural network ensemble
(b) Evidential deep learning
Figure 8: Results for different hyperparameter settings for neural network ensemble and evidential deep learning. The columns contain the results for learning rates 0.0001, 0.001 and 0.01. The color of the bars indicates the size and depth of the hidden layers of the network. Small, medium and large has 1, 3, and 8 hidden layers with 20, 200 and 2000 nodes respectively. The first and second rows of plots shows the sample mean and sample standard deviation
of the estimated class probability for the out-of-distribution test points. The bottom two rows of plots shows the sample mean and sample standard deviation of the estimated uncertainty for the out-of-distribution test points. The error bars indicate the 2.5-97.5 percentiles.
As an additional test for OOD performance we created a dataset with two statistically identical classes by using the parameters of class 1 in dataset A when generating data for both classes. We ran the same hyperparameter grid search for all deep learning models. The results for NNE and EDL are presented in Figure 8(b). NNE, CL and MCD all produced OOD class probability estimates which varied greatly both across hyperparameter values and as a function of polar angle. EDL stood out as the only algorithm, which (for small enough $\lambda_{0}$) produced estimated probabilities of 0.5. Interestingly, the uncertainty was still close to zero, which means it can not be used to detect OOD data and could lead to overconfident estimates. The GP and DPMM algorithms struggled with this dataset, especially for large numbers of data points so we could only evaluate performance on 2000 or less training data points. The GP hyperparameter optimization produced an output variance, which quickly approached zero as more data was added, leading to close to zero uncertainty everywhere, while the estimated probability stayed close to 0.5. For in distribution data, the DPMM gave probability estimates which were pretty noisy, as expected, but fluctuate around 0.5. Estimates for OOD data were approximately 0.5 and uncertainties were close to $0.5/\sqrt{3}$, in agreement with the chosen prior.
@@ 6 Discussion and conclusion
  • the average uncertainty on the test set go down as we increase the number of training data points for all algorithms except evidential deep learning which did not exhibit this behavior under the hyperparameters we tested, and its uncertainties seem hard to interpret in the case of our data;
@@
  • the out-of-distribution average uncertainty as a function of number of training data points is high and goes down for the Gaussian process classifier and Dirichlet process mixture model, and is low or zero and constant for the fully connected neural network models with ReLU activation functions;
@@
  • for out-of-distribution data all fully connected neural network models with ReLU activation functions estimated extreme probabilities, while the nonparametric Bayesian models estimate probabilities around 0.5, with the Dirichlet process mixture model estimating slightly higher probabilities as the number of training data points increase, indicating some numerical instability or other error in implementation.
@@
The DPMM is explicitly defined as having a $\text{Beta}(1,1)$ prior for class probability, and results support the qualitative correctness of the approximate posterior. For the neural networks and GP, the prior is implicitly defined, and so it is harder to evaluate the cohesiveness of the results. The approximate posterior distribution of the deep learning models behaves as if the probability mass is concentrated at 0 or 1. It is tempting to interpret this as an approximate posterior distribution with $\text{Beta}(0,0)$ priors for class probability. This interpretation suggests that changing the implicit prior for the deep learning models through modifying the loss function could change the observed behavior. However, it may not be a good idea to assign those kind of interpretations post-hoc, as we might put too much trust in a method without a principled theoretical foundation. Instead, we might consider if the approximations introduced by the approximate Bayesian inference methods (ensembles, MCD, variational approaches, etc.) are the cause of the deviation from desired results.
@@
Although neural networks with rectified linear (ReLU) activation functions can in theory approximate any nonlinear function of the data locally, they eventually become simple linear functions of the input features if one moves in any direction far enough (Hein et al., 2019). This property of piecewise linear neural networks, combined with using the softmax function for normalization of the last layer of the network, gives rise to class probabilities that asymptotically approach extreme values of zero or one. Therefore the estimated extreme probabilities for out-of-distribution data may simply be a feature of using ReLU neural networks as the family of parameterized test distributions in approximate inference. Similar to more well known parametric methods such as linear regression we should not expect our model to be informative or correct if we extrapolate too far. Changing the architecture and activation functions to better reflect the kind of statistical model we seek is an obvious step to mitigate the problem of model misspecification. In this study we tried changing the activation function to one that is bounded to see if this would help the situation, but our results showed that this was not a robust solution although for some models it seemed to improve OOD uncertainties (see Appendix C).
@@
This study only investigates data which is sampled from a relatively simple two dimensional distribution and therefore does not represent the complexities present in real-world data. It is noteworthy that even for this relatively simple problem, many of the learning algorithms do not produce uncertainties which can be used by scientists to decide if they should trust the class probability estimates. Although the uncertainty estimates produced by the deep learning algorithms presented in this study were not especially useful for this type of data, they might work better for data with for example discrete or bounded distributions, or for data with statistically independent input features. We strongly recommended that scientists create toy data that more closely matches the mathematical properties of their own data of interest to see how the different machine learning algorithms can be expected to behave in their case.
Regardless of the mathematical reason for the observed behavior and the inherent limitation of empirical studies, we can take away some key lessons from the experiment. The studied methods for class probability and uncertainty estimates do not necessarily exhibit the behavior we expect or desire for out-of-distribution data. This may be partly due to a lack of understanding of what can reasonably be expected from the algorithms in question, like the fact that it is not possible to have unbiased frequency estimates which also extrapolate to a specific value in the $N\rightarrow 0$ limit. It can also be because we do not yet have a clear understanding of the implicit statistical models created by our algorithms or the inherent limitations of the approximations we make in our inference algorithms to facilitate faster learning and compute intractable integrals. In any case, this study should serve as a reminder that uncertainty quantification for machine learning is still a largely unsolved problem. Sanity checks with intuitive toy-problems such as this one are a useful tool to explore the way these algorithms behave, and we hope this study can serve as an inspiration and a guide for future machine learning development work in the natural sciences.